OpenStudy (anonymous):

: You have an audio file of a 8.00 minute song. The song is sampled at 60.0 KHz with each sample taking 16 bits. What is the size of the audio file? Give your answer in Kbytes. , , @Physics

OpenStudy (stormfire1):

8 minutes = 8 * 60s = 480s 60 khz = 60,000 samples per second Your total samples are 480 * 60000 = 28800000 Each sample takes 2 bytes (16 bits) so you end up with 57,600 kbytes

OpenStudy (anonymous):

oh lol thanks, for the samples per sec i used the lossless compression rate, so I used 2*60k = 120k samples per sec. but what you have is the answer according to my homework console. :D

OpenStudy (stormfire1):

Well I geuss that's good. I've never been asked to calculate it before :)

OpenStudy (stormfire1):

Lossless or not, you should still be using 60,000 since that's all the samples you're taking every second

OpenStudy (anonymous):

hahah props to you skills :D i would not know how to answer this on the first time i see it. but there was a question too that asked for how many samples per second should be taken for lossless compression, and the right way was to multiply the frequency by 2 and that should be the samples taken per second, so yeah.. :D

OpenStudy (stormfire1):

Ok, can see that for that question :)

OpenStudy (stormfire1):

err *I can see*...

OpenStudy (anonymous):

haha yeah. thanks again!! :D

OpenStudy (stormfire1):

n p :)

OpenStudy (anonymous):

Use dimensional analysis. Each fraction is equivalent to 1, because they are the relevant conversion factors for the given units. The units all cancel and you are left with kb. \[8\min \times \frac{60s}{1\min} \times \frac{60,000samples}{1s} \times \frac{2bytes}{1sample} \times \frac{1kb}{1024bytes}\]

OpenStudy (stormfire1):

Good point..I should have done that. If I had, I would have realized that I needed 1024 vs 1000 bytes per kb :) So my answer is slightly off

OpenStudy (stormfire1):

For the record, it should have been 56,250 kb